Use the method of variation of parameters to find a particular solution of the following differential equation. y - by +9y
Answer #1

30 4-6y+90 - piesa noua au villany equation suc h en m?: 6m +9 30. (m-3) (m-3 2-0.. m = 3,3 complementary Solution is cf =to (4 ч.) - 9. : 32 г мез - зах - я кзе*, * 3 g - x3e + е за со (915) - Ге 37 320 . xe te 2 - e г 39 No ( 7 2 4c2 3x 3 xV: [Ч, СУ) dч - РеЎ? » (Я •42) х = . Г р() - ау, 1 ч 2 та — — ве 3 + Х аў у е = x e – m° 3 Уptx) = xe3x — : 6 90 Thu x w(),